La bibliographie de la partie ”Tomographie”

A. Coulais

4 avril 2003

Références

HER79_Bra [1] R. Bracewell, Image reconstruction in Radio Astronomy, in Image reconstruction from projections, G. Herman, ed., vol. 32 of Topics in applied physics, Springer Verlag, 1979, pp. 81–104.

Brem95 [2] M. Bremer, A 3D iterative deprojection technique. I. Development of the algorithm and tests, A&A Supplement Series, 112(0):551–566, mois inconnu 1995, .

HER79_BGH [3] T. Budinger, G. Gullberg, and R. Huesman, Emission computed tomography, in Image reconstruction from projections, G. Herman, ed., vol. 32 of Topics in applied physics, Springer Verlag, 1979, pp. 147–246.

Davi94 [4] J. Davila, Solar tomography, The Astrophysical Journal, 423(0):871–877, Mars 1994, .

DaTh92 [5] J. Davila and W. Thompson, A rotating tomographic imager for solar extreme–ultraviolet/soft X–ray emission, The Astrophysical Journal, 389(0):L91–L93, Avril 1992, .

DaGr79 [6] M. Davison and F. Grünbaum, Convolution algorithms for arbitrary projections angles, IEEE Trans. on Nuclear Sciences, 26(2):2670–2673, Avril 1979, .

DeBr95 [7] A. Delaney and Y. Bresler, Multiresolution tomographic reconstruction using wavelets, IEEE Trans. of image processing, 4(6):799–813, Juin 1995, .

GoHe74 [8] R. Gordon and G. Herman, Three–dimensional reconstruction from projections: a review of algorithms, International Journal of Cytology, 38:111–151, 1974, .

Her79 [9] G. Herman, ed., Image reconstruction from projections, vol. 32 of Topics in applied physics, Springer Verlag, 1979.

HGGB77 [10] R. Huesman, G. Gullberg, W. Greenberg, and T. Budinger, RECLBL library users manual, Lawrence Berkeley laboratory, University of California, 1977.

Kak84 [11] A. Kak, Image reconstruction from projections, in Digital Image Processing Techniques, M. Ekstrom, ed., vol. 2 of Computational Techniques, Academic Press, 1984, livre 4, pp. 111–170.

Kak85 [12] A. Kak, Tomographic imaging with diffracting an nondiffracting sources, in Array signal processing, J. Justice, N. Owsley, J. Yen, and A. KAK, eds., Prentice-Hall, 1985, livre 6, pp. 351–428.

Lea91 [13] D. Leahy, Deprojection of emission in axially symmetric transparent systems, A&A, 247(0):584–589, mois inconnu 1991, .

LoNa83 [14] A. Louis and F. Natterer, Mathematical problems of computerized tomography, Proc. IEEE, Special issue on Tomography, 71(3):379–389, Mars 1983, .

Mly93 [15] J. Mlynář, First computer tomography experimental results on the CASTOR tokamak, in Proceedings of contribution papers, F-2 section, physics of plasmas and ionized media, Faculty of mathematics and physics, Charles University, Prague, Septembre 1993, pp. 177–182.

Nil82 [16] R. Niland, Maximum reliable information obtainable by tomography, Journal of the Optical Society of America, 72(12):1677–1682, Décembre 1982, .

Rob84_Nil [17] R. Niland, Maximum reliable information in limited angle tomography, in Indirect Imaging, J. Roberts, ed., URSI, Australie 1983, 1984, pp. 325–331.

PGM96 [18] F. Peyrin, L. Garnero, and I. Magnin, Introduction à l’imagerie tomographique 2D et 3D reposant sur une propagation en ligne droite. Cas de la tomographie par rayon X par émission et par ultrasons, Traitement du signal, 13(4):380–450, Septembre 1996, .

PrWi90 [19] J. Prince and A. Willsky, Constrained sinogram restoration for limited-angle tomography, Optical Engineering, 29(5):535–544, Mai 1990, .

PrWi92 [20] J. Prince and A. Willsky, Hierarchical reconstruction unsing geometry and sinogram restoration, soumis à IEEE Trans. on image processing, 0, Février 1992, .

HER79_Row [21] R. Rowland, Computer implementation of image reconstruction formulas, in Image reconstruction from projections, G. Herman, ed., vol. 32 of Topics in applied physics, Springer Verlag, 1979, pp. 9–80.

SNLIH81 [22] T. Sato, S. Norton, M. Linzer, O. Ikeda, and M. Hirama, Tomographic image reconstruction from limited projections using iterative revisions in image and transform spaces, Applied Optics, 20(3):395–399, Février 1981, .

SwBa77 [23] W. Swindell and H. Barrett, Computerize tomography: taking sectional X rays, Physics Today, 0(0):32–41, Décembre 1977, .

UnAl96 [24] M. Unser and A. Aldroubi, A review of wavelets in biomedical applications, Proc. IEEE, Special issue on Wavelets, 84(4):626–638, Avril 1996, .